OpenStudy (anonymous):

4/x + 6/x+9 = (* / they are fractions)

OpenStudy (anonymous):

solve g^x+2 = (1/27) ^x=12

OpenStudy (anonymous):

log2(72)-2log2(3)

OpenStudy (anonymous):

4/x + 6/x+9 10x +36 / x(x + 9)

OpenStudy (anonymous):

evaluate a) 16^1.25 b) 125^4/3 c) 8^2/3 x 4^5/2

OpenStudy (anonymous):

a) 16^1.25 16^1 1/4 16^5/4 2^5 = 32

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

i dont understand how you got that

OpenStudy (anonymous):

u agree that .25 = 1/4 ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[16^{1.25}=16^{5/4}\] does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(2^{4})^{\frac{ 5 }{ 4 }}\] understand? cause 16 = 2^4

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

do b and ill correct if for u :) cool?

OpenStudy (anonymous):

125^4/3 (5^4)^1.25 this is all i can get?

OpenStudy (anonymous):

125 = 5^3 try it now :)

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

\[125^{4/3}\]\[(5^{3})^{\frac{ 4 }{ 3 }}\] better?